The function f(x) = 1 + 1.5ln(x+1) models the average number of free-throws a basketball player can make consecutively during practice as a function of time, where x is the number of consecutive days the basketball player has practiced for two hours. After 54 days of practice, what is the average number of consecutive free-throws the basketball player makes?
Answer by ankor@dixie-net.com(22740)

The function f(x) = 1 + 1.5ln(x+1) models the average number of free-throws a basketball player can make consecutively during practice as a function of time, where x is the number of consecutive days the basketball player has practiced for two hours. After 54 days of practice, what is the average number of consecutive free-throws the basketball player makes?
:
f(x) = 1 + 1.5*ln(x+1)
:
Substitute 54 for x in the equation:
f(54) = 1 + 1.5*ln(54+1)
f(54) = 1 + 1.5*ln(55)
f(54) = 1 + 1.5 * 4.00
f(54) = 1 + 6
f(54) = 7 free throws in a row